OpenStudy (anonymous):

Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find a if b = 6 and c = 10. A) a = 8 B) a = 4 C) a = 64 D) a = 16

Parth (parthkohli):

Use Pythagorean. \(\Large \color{Black}{\Rightarrow 6^2 + b^2 = 10^2 }\) \(\Large \color{Black}{\Rightarrow 36 + b^2 = 100 }\) \(\Large \color{Black}{\Rightarrow b^2 = 64 }\) \(\Large \color{Black}{\Rightarrow b = \sqrt{64} = 8 }\)

OpenStudy (anonymous):

Then how would you do ... Say that a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. Find c if a = 15 and b = 20. A) c = 13 B) c = 25 C) c = 35 D) c = 625